OpenStudy (anonymous):

What is the slope-intercept form of this equation? 3x + 9y = 18 A. y = -1/3x + 6 B. y = -1/3x + 2 C. y = 1/3x + 6 D. y = 2/3x + 2

OpenStudy (campbell_st):

subtract 3x from both sides of the equation. then divide both sides by 9. this will put the equation in slope intercept form.

OpenStudy (anonymous):

thats the thing im horrible at slopes and it confuses me

OpenStudy (anonymous):

@dan815 @campbell_st

OpenStudy (campbell_st):

3x + 9y = 18 -3x -3x ------------------- 9y = -3x + 18 now divide every term by 9

OpenStudy (anonymous):

9 / 9 =1 -3 / 9 = -0.333333333333333 @campbell_st

OpenStudy (campbell_st):

can you simplify -3/9 and leave it as a fraction

OpenStudy (campbell_st):

and if you divide the constant by 9 the answer should be obvious

OpenStudy (anonymous):

i'm so confused )':